Everyday to avoid traffic driver morning commute, Yunru Takes the back roads from Her home to her workplace. There is less traffic during the afternoon commute home from her workplace so she always takes the highway instead of the backroads. Taking the back roads is 7 miles longer than taking the highway, but her total commute distance is 57 miles long. How many miles long is her morning commute?
A) 25
B)32
C)39
D)43

Answer by ikleyn(52792)

.
Let x = the length of the morning commute and Y = the length of the back road, im miles.


Then from the condition, you have these two equations


    Y + X = 57     ( her total commute distance )


    Y - X =  7     ( back road is 7 miles longer than . . . )


Subtract the second equation from the first one. You will get


    2X = 50;  hence  X = 50/2 = 25 miles.


It is your answer.